The amount of money that a salesperson makes varies directly as the total amount of sales made. If the salesperson receives $7,500 for selling a house for $120,000, how much will he or she make if he or she sells a house for $150,000? Set up a proportion. I figured out the problem but am not sure how to show a check on it. Please help.
Answer by mananth(16946)

7500/120,000 = x/150,000
7500*150000/120000=x
x= $9375